# Improve the Precision of Area Under the Curve Estimation for Recurrent Events Through Covariate Adjustment

### v2

@@ 3.1 Data Generation
The simulation studies aim to assess the accuracy of our variance formula and compare the statistical power between the AUC and LWYY methods. Three independent baseline covariates $\boldsymbol{X}^{*}=(X_{1}^{*},X_{2}^{*},X_{3}^{*})$ are considered. Specifically, $X_{1}^{*}$ follows a Bernoulli distribution with probability of 0.5 while $X_{2}^{*}$ and $X_{3}^{*}$ independently follow a normal distribution with mean of 0 and standard deviation of 2.
@@
There are a total of 2000 subjects. For each subject $i$, the death time $D_{i}=5/365+\omega_{i}$, where $\omega_{i}$ follows an exponential distribution with the mean as $\left\{0.05\hbox{exp}(X_{i}^{*T}\xi)\right\}^{-1}$, where $\xi=(0.1,0.1,0.1)^{T}$. The maximum follow-up time for each subject is 2 years. The independent censoring $C_{i}$ is uniformly distributed on $(1,2)$.
@@
  • Case 2: Recurrent events are generated using the conditional intensity function: $\rho^{(2)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(1)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(1)}(t)=-0.25\theta t^{2}+\theta t$. The treatment effect is characterized by a time-varying $\theta^{(1)}(t)$, which moves from 0 at $t=0$ to $\theta$ at $t=2$.
@@
  • Case 3: Recurrent events are generated from the conditional intensity function: $\rho^{(3)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(2)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(2)}(t)=-\theta t^{2}+2\theta t$. $\theta(t)$ moves from 0 at $t=0$, reaches a peak of $\theta$ at $t=1$, and backs to 0 at $t=2$.
@@
  • Case 4: Recurrent events are generated from the conditional intensity function: $\rho^{(4)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(3)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(3)}(t)=-0.25\theta t^{2}+\theta$. $\theta(t)$ moves from $\theta$ at $t=0$ to 0 at $t=2$.
@@
  • Case 5: Let $\Delta V_{ij}$ be the gap time between the $(j-1)$th and $j$th recurrent event for subject $i$. Define $\Delta V_{ij}=\hbox{exp}(-\theta I_{i}+X_{i}^{*T}\eta-0.7)+\epsilon_{j}$, where $\epsilon_{j}$ is an independent random variable following an exponential distribution with mean as 0.25. The ordered $j$th recurrent event time for subject $i$ is then $\sum_{k=1}^{j}\Delta V_{ik}$, for $\sum_{k=1}^{j}\Delta V_{ik}<T_{i}$, with the 00th recurrent event time being 0.
The LWYY model is correctly specified in Case 1 but misspecified in Cases 2 through 5. In Cases 2 through 4, the treatment effect varies over time: Case 2 features a late treatment effect scenario, where the treatment effect gradually increases and reaches its full extent by the end; Case 3 exhibits a peak treatment effect scenario, with the effect being strongest at a specific point and diminishing before and after that peak; and Case 4 presents an early treatment effect scenario, where the effect is immediate at the beginning and then diminishes over time. In these cases, the constant treatment effect parameter estimated by the LWYY model does not accurately reflect the varying treatment effects, and is hard to interpret. In Case 5, the treatment affects recurrent events by reducing the gap time between them. The LWYY model is poorly specified in this case, and the estimated treatment effect may be thus misleading.
3.2 Simulation Results
@@
The mean follow-up across the five cases and two randomization schemes is 1.44 years, with 8% of subjects experiencing mortality during the study period. The average number of recurrent events per subject is the same between the two randomization schemes but varies slightly between cases. For the placebo arm, the average number of recurrent events per subject is 0.68 for Cases 1 to 4, and 0.81 for Case 5. For the treatment arm, with $\theta=-0.32$, the numbers are 0.49, 0.51, 0.55, 0.58, and 0.72 for Cases 1 to 5, respectively.
Figure 1 shows the MCFs for the five cases with the simple randomization scheme, using a total sample size of 100,000 with equal allocation to generate smoother curves. $\theta$ is set to -0.32. The MCF at $t=2$ is greater than the corresponding average number of events per subject because the MCF implicitly imputes events for those who have been censored or died, using the event rate estimated from subjects remaining at risk, modified by the corresponding survival probability, as suggested by Equation 1.
@@
The treatment effect is constant in Case 1 but not in Cases 2 to 5. However, it is difficult to detect the time-varying treatment effects from MCFs. To better view the treatment effect over time, the black curve in Figure 1 represents $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ over time. Despite some instabilities at earlier times, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ is constant over time in Case 1. In Case 2, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ decreases over time. In Case 3, the trend of $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ is not very clear before year 1, but after year 1, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ shows an obvious upward trend. In Case 4, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ increases over time, and in Case 5, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ remarkably increases early on and becomes stable near the tail.
3.2.1 Adjusted AUC vs Unadjusted AUC
Figure 2 shows the distribution of estimated standard errors $\sqrt{\hat{\sigma}_{L,R}^{2}/n}$ (unadjusted) and $\sqrt{\hat{\sigma}_{CL,R}^{2}/n}$ (adjusted) from 5000 simulations under different randomization schemes and cases, with $\theta=-0.32$ and $\tau=2$. Similar results for $\sqrt{\hat{\sigma}_{L,\Delta}^{2}/n}$ and $\sqrt{\hat{\sigma}_{CL,\Delta}^{2}/n}$ are presented in the Section C.1 of the Supplementary Materials. Yellow diamonds represent the Monte Carlo standard deviations of 5000 $\hat{\mathcal{U}}_{L,R}$ (unadjusted) and $\hat{\mathcal{U}}_{CL,R}$ (adjusted).
In the unadjusted analysis, the yellow diamond for simple randomization is higher compared to SPB randomization. This occurs because SPB randomization reduces variability by enforcing balance with respect to $X^{*}_{1}$ and $X^{*}_{2}$.
@@ 3.2.1 Adjusted AUC vs Unadjusted AUC
For simple randomization in the unadjusted analysis, as well as for both simple and SPB randomization in the adjusted analysis, the boxplot bodies are thin and includes the yellow diamond, despite some outliers. This empirically verifies the asymptotic properties outlined in Theorem 2, suggesting that our variance formulas are accurate and that the covariate adjustment approach is applicable to both simple and covariate-adaptive randomization. Conversely, in SPB randomization, the boxplots from the unadjusted analysis are positioned above the corresponding yellow diamonds, indicating that the estimated standard errors overestimate the actual standard deviations, making the analyses conservative. This observation is consistent with the FDA guidance on covariate adjustment [Food and Drug Administration (2023), FDA] and further supports the notion that standard errors can be overestimated when covariate-adaptive randomization schemes are ignored. In practice, if baseline covariates are balanced during randomization, these covariates should also be adjusted for in the analysis.
@@
Figure 3 compares the power of unadjusted AUC with adjusted AUC based on 5000 simulations. Tests on the difference and ratio of AUCs show identical power. The figure demonstrates that the adjusted analysis consistently yields higher power across various cases and randomization schemes. This empirically supports the efficiency gain suggested by Theorems 1 and 2. Additionally, with SPB randomization, the Type-I errors for the unadjusted analysis are below the nominal level. This suggests that the unadjusted analysis that does not account for covariate-adaptive randomization is conservative.
3.2.2 AUC vs LWYY
One may be interested in the relative power between LWYY and AUC, although this is not the primary focus of this paper. Figure 3 also compares the power of AUC (both unadjusted and adjusted) with that of LWYY (both unconditional and conditional), acknowledging that the null hypotheses for AUC and LWYY differ, and that adjusted LWYY provides conditional estimates. The AUC is calculated over the interval $(0,\tau)$, whereas the LWYY is not limited to this interval and evaluates differences across the entire follow-up period. The unconditional LWYY estimate is derived from the model $\mu_{1}(t)=\mu_{0}(t)\hbox{exp}(\theta)$, where $\mu_{1}(t)$ and $\mu_{0}(t)$ are the MCFs of the treatment and placebo arm, respectively. The conditional LWYY estimate is derived from the model $\mu(t\mid\boldsymbol{X}^{*})=\mu_{0}(t)\hbox{exp}(\theta j+\boldsymbol{X}^{*T}\eta)$ for $j=0,1$.
The relative power between AUC and LWYY is similar across different randomization schemes but varies across cases. We focus on the comparison between conditional LWYY and adjusted AUC. In Case 1, where the LWYY model is correctly specified, the LWYY model has higher power because its inference is based on likelihood, providing optimal power. In Case 2, where the treatment effect is larger at later times, although the LWYY model is misspecified, it still has better power than AUC. In Case 3, the adjusted AUC has comparable power to the conditional LWYY, but the unadjusted AUC has slightly better power than the unconditional LWYY. In Cases 4 and 5, where the treatment effect is larger at earlier times, AUC has better power than LWYY.
These results suggest that AUC may provide better power than LWYY if the treatment effect decreases over time. This is because the test statistic of AUC assigns lesser weights to later differences compared to the LWYY model [Tian et al., 2018, Uno and Horiguchi, 2023]. Therefore, when the treatment effect is larger at later study time points, AUC tends to be less powerful than LWYY. Conversely, if the treatment effect is smaller at later study time points, AUC is likely to be more powerful than LWYY. In the FINEARTS-HF study, the treatment effects on mean cumulative events appear to diminish over time, resembling Case 4 in our simulations, where AUC might demonstrate greater power than LWYY.
From an interpretative perspective, in Cases 2 to 5, where the treatment effect varies over time, the AUC provides a more clinically meaningful interpretation, as the LWYY model erroneously assumes and estimates a constant treatment effect. Furthermore, the adjusted AUC preserves the unconditional interpretation, while the adjusted LWYY targets a conditional estimand.
The choice of $\tau$ may change the relative power between LWYY and AUC. With a smaller $\tau$, AUC counts a smaller number of recurrent events, which might suggest diminished power for AUC. One might then question whether AUC can still have better power than LWYY with a smaller $\tau$. Interestingly, Table S1 in the Section C.2 of the Supplementary Materials shows that in Cases 4 and 5, AUC even has better power with $\tau=1.7$ than with $\tau=2$. This is because, in these cases, setting a smaller $\tau$ ignores the time period when the treatment effect is relatively smaller, so $\hat{U}_{1}(1.7)/\hat{U}_{0}(1.7)<\hat{U}_{1}(2)/\hat{U}_{0}(2)$, as shown in Figure 1. These results suggest that in cases where the treatment effect decreases over time, AUC with a smaller $\tau$ is likely to yield even higher power.
@@ 4.1 HF-ACTION
In the analysis, we use an age indicator (whether a patient is over 60) as a covariate. Table 1 shows the estimated ratio and difference of AUCs between the two arms over the time window $[0,4]$. After adjusting for age, the standard error of the log ratio of AUCs decreases from 0.123 to 0.121 (a 1.6% reduction), and the standard error of the difference in AUCs decreases from 0.877 to 0.868 (a 1% reduction). These reductions suggest that covariate adjustment improves the precision of our estimates.
@@ 4.2 COPERNICUS
The lower panel of Figure 4 shows the MCF for both arms, along with the ratio of AUCs, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$, over time (black curve). The treatment effect appears to remain relatively constant over time, making the LWYY model an appropriate choice for this data. However, for illustration purposes, we will proceed with an AUC analysis. Covariates in this analysis include sitting heart rate, history of myocardial infarction, presence of concomitant disease, diabetes, and the use of diuretics, ACE inhibitors, digitalis, amiodarone, and anticoagulants. Table 1 provides the estimated ratio and difference of AUCs between the two arms over the time window $[0,2.2]$. After adjusting for covariates, the standard error of the log ratio of AUCs decreases from 0.074 to 0.072 (a 2.7% reduction), and the standard error of the difference in AUCs decreases from 0.124 to 0.121 (a 2.4% reduction). These reductions further demonstrate that covariate adjustment improves the precision of our estimates.
@@ 5 Discussion
Theoretical asymptotic distributions demonstrate that our approach achieves efficiency gains, and extensive simulations across various scenarios and real data analysis confirm its effectiveness in improving statistical power. This nonparametric approach offers several advantages. First, this approach is model-free. Although we use simple linear regression to represent $\mathbb{E}(\hat{P}_{i}^{\Delta})$, $\mathbb{E}(\hat{Q}_{i}^{\Delta})$, $\mathbb{E}(\hat{P}_{i}^{R})$, and $\mathbb{E}(\hat{Q}_{i}^{R})$ during adjustment, the approach is valid regardless of whether the linear regression model correctly specifies the true expected values. In other words, its validity does not depend on parametric modeling assumptions [Zhang et al., 2008]. The choice of linear regression is for implementation convenience and offers good theoretical properties. This model-free characteristic aligns with the “minimal statistical assumption" principle advocated in the recent FDA guidance on covariate adjustment [Food and Drug Administration (2023), FDA]. Second, unlike conventional regression-based approaches for covariate adjustment, which can lose power if the model is misspecified, the nonparametric covariate adjustment guarantees an efficiency gain over the unadjusted test. Third, this approach is universally applicable to both simple and covariate-adaptive randomization schemes, unlike some nonparametric methods that are limited to simple randomization [Lu and Tsiatis, 2008]. Lastly, unlike non-linear regression models where the distinction between conditional and unconditional estimates is pronounced (known as non-collapsibility), this approach does not alter the unconditional interpretations.
@@
We also compared the power of the LWYY model with the AUC method, acknowledging their different null hypotheses. As the AUC method tends to assign lesser weights to later study times, it tends to have higher power when the treatment effects are large at earlier times. This observation aligns with previous studies that compared the power of restricted mean survival time (RMST) and the Cox model in time-to-event settings [Tian et al., 2018]. In such cases, specifying a smaller $\tau$ in AUC does not necessarily reduce power, even though events near the tail are excluded from the analysis.
@@
To the best of our knowledge, this is the first paper to provide a robust, unconditional covariate-adjusted inference for area under the curve estimand in the recurrent event setting, covering covariate-adaptive randomizations. The curve can represent not only the mean cumulative function but also the survival curve. The application of this approach to RMST is straightforward and discussed in the Section D of the Supplementary Materials. Future research on AUC may include sample size calculation and interim analysis.

### v3

@@ 3.1 Data Generation
Three independent baseline covariates $\boldsymbol{X}^{*}=(X_{1}^{*},X_{2}^{*},X_{3}^{*})$ are considered. Specifically, $X_{1}^{*}$ follows a Bernoulli distribution with probability of 0.5 while $X_{2}^{*}$ and $X_{3}^{*}$ independently follow a normal distribution with mean of 0 and standard deviation of 2.
@@
There are a total of 2,000 subjects. For each subject $i$, the death time $D_{i}=5/365+\omega_{i}$, where $\omega_{i}$ follows an exponential distribution with the mean as $\left\{0.05\hbox{exp}(X_{i}^{*T}\xi)\right\}^{-1}$, where $\xi=(0.1,0.1,0.1)^{T}$. The maximum follow-up time for each subject is 2 years. The independent censoring $C_{i}$ is uniformly distributed on $(1,2)$.
@@
  • Case 2: Recurrent events are generated using the conditional intensity function: $\rho^{(2)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(1)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(1)}(t)=-0.25\theta t^{2}+\theta t$. The treatment effect is characterized by a time-varying $\theta^{(1)}(t)$, which moves from 0 at $t=0$ to $\theta$ at $t=2$. This case features a late treatment effect scenario, where the treatment effect gradually increases and reaches its full extent by the end.
@@
  • Case 3: Recurrent events are generated from the conditional intensity function: $\rho^{(3)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(2)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(2)}(t)=-\theta t^{2}+2\theta t$. $\theta(t)$ moves from 0 at $t=0$, reaches a peak of $\theta$ at $t=1$, and backs to 0 at $t=2$. This case exhibits a peak treatment effect scenario, with the effect being strongest at a specific point and weaker before and after that peak.
@@
  • Case 4: Recurrent events are generated from the conditional intensity function: $\rho^{(4)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(3)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(3)}(t)=-0.25\theta t^{2}+\theta$. $\theta(t)$ moves from $\theta$ at $t=0$ to 0 at $t=2$. This case presents an early treatment effect scenario, where the effect is immediate at the beginning and then diminishes over time.
@@
  • Case 5: Let $\Delta V_{ij}$ be the gap time between the $(j-1)$th and $j$th recurrent event for subject $i$. Define $\Delta V_{ij}=\hbox{exp}(-\theta I_{i}+X_{i}^{*T}\eta-0.7)+\epsilon_{j}$, where $\epsilon_{j}$ is an independent random variable following an exponential distribution with mean as 0.25. The ordered $j$th recurrent event time for subject $i$ is then $\sum_{k=1}^{j}\Delta V_{ik}$, for $\sum_{k=1}^{j}\Delta V_{ik}<T_{i}$, with the 00th recurrent event time being 0. Case 5 differs from the preceding cases in that the treatment affects recurrent events by reducing the gap time between them.
3.2 Results
@@
The mean follow-up across the five cases and two randomization schemes is 1.44 years, with 7.8% of subjects experiencing mortality during the study period. The average number of recurrent events per subject is the same between the two randomization schemes but varies slightly between cases. For the placebo arm, the average number of recurrent events per subject is 0.68 for Cases 1 through 4 and 0.81 for Case 5. In the treatment arm, with $\theta=-0.16$, the corresponding averages are 0.58, 0.59, 0.61, 0.63, and 0.72 for Cases 1 through 5, respectively. When $\theta=-0.32$, the averages decrease further to 0.49, 0.51, 0.55, 0.58, and 0.62 for Cases 1 through 5, respectively.
@@
Figure 1 shows the MCFs for the five cases with the simple randomization scheme, using a total sample size of 100,000 with equal allocation to generate smoother curves. $\theta$ is set to -0.32. The treatment effect is constant in Case 1 but not in Cases 2 to 5. However, it is difficult to detect the time-varying treatment effects from MCFs. To better view the treatment effect over time, the black curve in Figure 1 represents $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ over time. Despite some instabilities at earlier times, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ is constant over time in Case 1. In Case 2, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ decreases over time. In Case 3, the trend of $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ is not very clear before year 1, but after year 1, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ shows an obvious upward trend. In Case 4, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ increases over time, and in Case 5, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ remarkably increases early on and becomes stable near the tail.
Table 2 presents simulation results for the difference and ratio of AUCs from both unadjusted and covariate-adjusted analyses across various scenarios and randomization schemes, based on 5,000 simulations. For practical purposes, we use the Monte Carlo mean of the 5,000 estimated unadjusted differences and ratios of AUCs—denoted as $\Delta\hat{U}(\tau)$ and $\hat{R}(\tau)$—as empirical approximations of the true values $\Delta U(\tau)$ and $R(\tau)$, and label them as “Est.” “Bias” refers to the difference between the Monte Carlo mean of the 5,000 adjusted estimates ($\Delta\hat{U}^{(a)}(\tau)$ and $\hat{R}^{(a)}(\tau)$) and the corresponding “Est.” As expected, the “Bias” is consistently zero across all scenarios, empirically confirming that the proposed covariate adjustment procedure preserves the estimand. “Mean” and “Median” denote the mean and median of the 5,000 estimated standard errors (SEs) for each point estimate—$\Delta\hat{U}(\tau)$, $\Delta\hat{U}^{(a)}(\tau)$, $\hat{R}(\tau)$, and $\hat{R}^{(a)}(\tau)$—while “MC” represents the Monte Carlo standard deviation of the corresponding estimates. Coverage probability (CP) refers to the likelihood that the 95% confidence interval (CI) contains the respective “Est.” For the adjusted analyses, “Mean,” “Median,” and “MC” are nearly identical across all settings, and the CP remains close to the nominal level. In contrast, under the SPB randomization scheme, the unadjusted analysis yields “Mean” and “Median” values consistently larger than “MC,” with CP exceeding the nominal level—suggesting that unadjusted analyses that ignore covariate-adaptive randomization may be conservative. This observation aligns with FDA guidance on covariate adjustment, which recommends adjusting for baseline covariates in the analysis if they were balanced during randomization [Food and Drug Administration (2023), FDA]. Additional results for a smaller sample size (total of 400 subjects) are provided in Section D.1 of the Supplementary Materials. The results are similar to those presented in Table 2.
@@
To visually demonstrate the validity of our variance formulas and highlight the conservativeness of the unadjusted analysis under covariate-adaptive randomization, Figure 2 displays the distribution of estimated standard errors $\sqrt{\hat{\sigma}_{L,R}^{2}/n}$ (unadjusted) and $\sqrt{\hat{\sigma}_{CL,R}^{2}/n}$ (adjusted) from 5,000 simulations under different randomization schemes and cases, with $\theta=-0.32$ and $\tau=2$. Similar results for $\sqrt{\hat{\sigma}_{L,\Delta}^{2}/n}$ and $\sqrt{\hat{\sigma}_{CL,\Delta}^{2}/n}$ are presented in Section D.2 of the Supplementary Materials. Yellow diamonds represent the Monte Carlo standard deviations of 5,000 $\hat{\mathcal{U}}_{L,R}$ (unadjusted) and $\hat{\mathcal{U}}_{CL,R}$ (adjusted). In the unadjusted analysis, the yellow diamond for simple randomization is higher compared to SPB randomization. This occurs because SPB randomization reduces variability by enforcing balance with respect to $X^{*}_{1}$ and $X^{*}_{2}$. For simple randomization in the unadjusted analysis, as well as for both simple and SPB randomization in the adjusted analysis, the boxplot bodies are thin and include the yellow diamond, despite some outliers. This empirically verifies the asymptotic properties outlined in Theorem 1, suggesting that our variance formulas are accurate and that the covariate adjustment approach is applicable to both simple and covariate-adaptive randomization. Conversely, in SPB randomization, the boxplots from the unadjusted analysis are positioned above the corresponding yellow diamonds, indicating that the estimated standard errors overestimate the actual standard deviations, making the analyses conservative. This observation further supports the notion that standard errors can be overestimated when covariate-adaptive randomization schemes are ignored.
@@
Figure 3 compares the power of unadjusted AUC with adjusted AUC based on 5,000 simulations. Tests on the difference and ratio of AUCs show identical power. The figure demonstrates that the adjusted analysis consistently yields higher power across various cases and randomization schemes. This empirically supports the efficiency gain suggested by Theorem 1.
@@ 4.1 HF-ACTION
In the analysis, we use an age indicator (whether a patient is over 60) as a covariate. Table 3 shows the estimated ratio and difference of AUCs between the two arms over the time window $[0,4]$. After adjusting for age, the variance of the log ratio of AUCs decreases from 0.0151 to 0.0147 (a 2.65% reduction), and the variance of the difference in AUCs decreases from 0.7695 to 0.7526 (a 1.02% reduction). These reductions suggest that covariate adjustment improves the precision of our estimates.
@@ 4.2 COPERNICUS
The lower panel of Figure 4 shows the MCF for both arms, along with the ratio of AUCs, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$, over time (black curve). The treatment effect appears to remain relatively constant over time, making the LWYY model an appropriate choice for this data. However, for illustration purposes, we will proceed with an AUC analysis. Covariates in this analysis include sitting heart rate, history of myocardial infarction, presence of concomitant disease, diabetes, and the use of diuretics, ACE inhibitors, digitalis, amiodarone, and anticoagulants. Table 3 provides the estimated ratio and difference of AUCs between the two arms over the time window $[0,2.2]$. After adjusting for covariates, the variance of the log ratio of AUCs decreases from 0.0055 to 0.0052 (a 5.77% reduction), and the variance of the difference in AUCs decreases from 0.0153 to 0.0146 (a 4.79% reduction). These reductions further demonstrate that covariate adjustment improves the precision of our estimates.
4.3 Practical Implications
In this section, we demonstrate our covariate adjustment approach for AUC using two real-world clinical trials: the HF-ACTION trial and the COPERNICUS trial. For HF-ACTION, we have access only to a subgroup of participants with a single available covariate. The COPERNICUS trial includes all subjects, but some baseline covariates have missing values. While imputation is a viable strategy, we limit our analysis to covariates with complete data for illustrative purposes. Incorporating additional covariates would likely yield even greater efficiency gains.
The relative efficiency between two estimators is defined as the ratio of their variances. In clinical trial design, this can also be interpreted as the ratio of sample sizes required to achieve the same statistical power. In the COPERNICUS example, we observe an approximate 5% reduction in variance, which can be viewed as a potential 5% reduction in required sample size. A similar efficiency gain was observed in previous covariate adjustment studies [Ye et al., 2022, 2023, 2024].
Such efficiency gains are particularly meaningful in the context of modern clinical trials. As standard care continues to improve and event rates decline, larger sample sizes are often necessary to observe a sufficient number of events. For instance, while COPERNICUS (1997–2000) and HF-ACTION (2003–2007) enrolled fewer participants, many contemporary cardiovascular trials now exceed 10,000 subjects. A 5% reduction in required sample size could correspond to 500 fewer subjects in a 10,000-person trial—offering substantial savings in time, cost, and patient burden.
@@ 5 Discussion
Theoretical asymptotic distributions demonstrate that our approach achieves efficiency gains, and extensive simulations across various scenarios and real data analysis confirm its effectiveness in improving statistical power. This nonparametric approach offers several advantages. First, this approach is model-free. Although we use simple linear regression to represent $\mathbb{E}(\hat{P}_{i}^{\Delta})$, $\mathbb{E}(\hat{Q}_{i}^{\Delta})$, $\mathbb{E}(\hat{P}_{i}^{R})$, and $\mathbb{E}(\hat{Q}_{i}^{R})$ during adjustment, the approach is valid regardless of whether the linear regression model correctly specifies the true expected values. In other words, its validity does not depend on parametric modeling assumptions [Zhang et al., 2008]. The choice of linear regression is for implementation convenience and offers good theoretical properties. This model-free characteristic aligns with the “minimal statistical assumption” principle advocated in the recent FDA guidance on covariate adjustment [Food and Drug Administration (2023), FDA]. Second, unlike conventional regression-based approaches for covariate adjustment, which can lose power if the model is misspecified, the nonparametric covariate adjustment guarantees an efficiency gain over the unadjusted test. Third, this approach is universally applicable to both simple and covariate-adaptive randomization schemes, unlike some nonparametric methods that are limited to simple randomization [Lu and Tsiatis, 2008]. Lastly, unlike non-linear regression models where the distinction between conditional and unconditional estimates is pronounced (known as non-collapsibility), this approach does not alter the unconditional interpretations.
@@
To the best of our knowledge, this is the first paper to provide a robust, unconditional covariate-adjusted inference for the area under the curve estimand in the recurrent event setting, covering covariate-adaptive randomizations. The curve can represent not only the mean cumulative function but also the survival curve. The application of this approach to RMST is straightforward and discussed in Section E of the Supplementary Materials. While not the primary focus of this paper, we also examine the relative power of the LWYY and AUC approaches across various scenarios in Section D.3 of the Supplementary Materials, acknowledging that they test different null hypotheses. Future research on AUC may include sample size calculation and interim analysis.